I am struggling with understanding the problem below. It has been nearly 50 years since I did anything with log functions and I guess I don't have it anymore. Anyways here is the equation I need solved -
X= 10*(sqrt(n/k))*(0.5+0.5*log(b))
Assume n=30
k=5
b=40
Is there somewhere I can find the log of any number whether it be 40, 75, 100 etc ???

Answer by Alan3354(69443)

X= 10*(sqrt(n/k))*(0.5+0.5*log(b))
Assume n=30
k=5
b=40
Is there somewhere I can find the log of any number whether it be 40, 75, 100 etc ???
----------------
X= 10*(sqrt(n/k))*(0.5+0.5*log(b))
X= 10*(sqrt(30/5))*(0.5+0.5*log(40))
x = 10*sqrt(6)*(0.5 + 0.5*log(40)
x = 5sqrt(6) + 5*sqrt(6)*log(40)
Now it's calculator approxmations
x =~ 12.24745 + 12.24745*1.60206
x =~ 31.8686
--------------
Use a calculator for logs
A Casio fx-260 is $8 at Office Depot
There's one on the computer,
Start, Programs, Accessories, Calculator. Select View - Scientific